OpenStudy (anonymous):

Mamie has a bag containing three red counters, two blue counters and four green counters.She takes out a counter without looking in the bag. What is the probability that she doesnt pick a red counter

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Miracrown

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

@aayushi.somani

OpenStudy (anonymous):

@perl

OpenStudy (ribhu):

answer is 2/3

OpenStudy (aayushi.somani):

the probablity of having no red ones will be= B+G/total no. counters so it comes out to be 6/9= 2/3

OpenStudy (ribhu):

it can also be 1-(probability a blue or green counter is picked) = 1-(3/9) = 6/9 = 2/3

OpenStudy (anonymous):

thanks @aayushi.somani and @ribhu -willing to help unlike some people.